Andrea’s Fashions sold merchandise for $203000 cash during the month of July. Returns that month totaled $3100. If the company’s gross profit rate is 40%, Andrea’s will report monthly net sales revenue and cost of goods sold of
Answer #1

Answers :

Net sales = $199,900

Cost of Goods Sold = $ 119,940

Net sales = Total Sales - Sales Returns = 203,000-3,100 = $199,900(Answer)

Cost of Goods Sold = Net sales *(1- Gross Profit Ratio ) =199,900-(1-0.40) =199,900*0.60= $119,940 (Answer)